OpenStudy (anonymous):

8x-6z=38 2x-5y+3z=5 x+10y-4z=8 plz help solve each system of equation

OpenStudy (texaschic101):

2x - 5y + 3z = 5 -->(2) x + 10y - 4z = 8 ----------------- 4x - 10y + 6z = 10 (result of multiplying by 2) x + 10y - 4z = 8 -----------------add 5x + 2z = 18 8x - 6z = 38 5x + 2z = 18 -->(3) ---------------- 8x - 6z = 38 15x + 6z = 54 -----------------add 23x = 92 x = 92/23 x = 4 now sub 4 into an equation where y is eliminated 8x - 6z = 38 8(4) - 6z = 38 32 - 6z = 38 -6z = 38 - 32 -6z = 6 z = 6/-6 z = -1 now sub 4 in for x and -1 in for z in one of the original equations 2x - 5y + 3z = 5 2(4) - 5y + 3(-1) = 5 8 - 5y - 3 = 5 5 - 5y = 5 -5y = 5 - 5 -5y = 0 y = 0 check by subbing in known answers... x + 10y - 4z = 8 4 + 10(0) - 4(-1) = 8 4 + 0 + 4 = 8 4 + 4 = 8 8 = 8 (correct) answer : x = 4, y = 0, and z = -1

OpenStudy (anonymous):

texas alot of steps lol

OpenStudy (texaschic101):

yes...it has a lot of steps :)